A force of 85.5 N in the horizontal direction is required to set a 25.4-kg box, initially at rest on a flat surface, in motion. After the box begins to move, a horizontal force of 55.5 N is required to maintain a constant speed.

(a) What is the coefficient of static friction between the box and the surface?


(b) What is the coefficient of kinetic friction between the box and the surface?
Answer #1

here,

the mass of box , m = 25.4 kg

a)

let the coefficient of static friction between the box and the surface be us

the force required to set the box in motion , ffs = us * m * g

85.5 = us * 25.4 * 9.81

solving for us

us = 0.343

the coefficient of static friction between the box and the surface is 0.343

b)

let the coefficient of kinetic friction between the box and the surface be uk

the force required to keep the box in motion , ffk = uk * m * g

55.5 = uk * 25.4 * 9.81

solving for uk

uk = 0.22

the coefficient of static friction between the box and the surface is 0.22